(1 point) Starting with a =-1, b = 1, do 4 terations of golden section search to estimate where f(x)-(r-sin()) reaches a mini7 significant digits please
Answer #1

Solution: Note that 5-1 = 0.6180340 R= is called the golden-ratio. It is the positive root of the quadratic equation R^2+R-1=0.

The given funcion is f()-sin() . The Golden Section Search Method for finding the minimum of the function f()-sin() in the interval [-1,1] is given in the following table. The necessary steps are

Step1:
n=0

If f(x_1^0)<f(x_2^0)\\ , then set a^1=x_2^0,b^1=b^0 .

Step 2:
rl R(b -a)

Check  f(r) < f(x)), then set a^2=a^1,b^2=x_2^1 .
Continuing in this way we can find the minimum.


All the computations reported here are rounded to seven decimal of places

a b c d f(c) f(d)
i=0 -1 1 0.236068 -.236068 -0.1781534 0.2896096
i=1 -0.236068 1 0.5278641 0.2360679 -0.2250488 -0.1781534
i=2 -0.236068 0.5278641 0.2360680 0.0557281 -.01781534 -0.05259364
i=3 0.2360680 0.5278641 0.4164079 0.3475242 -0.2310824 -0.2197980
i=4 0.4164079 0.5278641 0.4852916 0.4589804 -0.2309584 -0.2323713